[figure: 9.7 cm]

ELECTRONIC AND TRANSPORT PROPERTIES OF Dy_{2}Co_{7}B_{3}

COMPOUND

A. Kowalczyk^{a}, G. Chełkowska^{b}, T Toliński^{a}, V. Ivanov^{c} aInstitute of Molecular Physics, PAS, Smoluchowskiego 17, 60-179 Poznań, Poland bInstitute of Physics, Silesian University, Uniwersytecka 4, 40-007 Katowice, Poland

cGeneral Institute of Physics, RAS, Vavilov 38, Moscow, Russia

New results of the electronic and transport properties (electrical resistivity) of Dy_{2}Co_{7}B_{3} compound are reported. This alloy crystallizes in the hexagonal Ce_{2}Co_{7}B_{3} type of structure. The electronic structure has been studied by the X-ray photoemission spectroscopy (XPS). The valence band of the XPS spectra is determined mainly by the Co(3d) and Dy(4f) bands. The peaks positions are in a good agreement with the binding energies of a metallic dysprosium. The resistivity at low temperatures shows a T^{2} dependence. At higher temperatures the resistivity is a nonlinear function of temperature, which indicates an electron-phonon interaction in the presence of a small s-d scattering.

[figure: 13.4 cm]

Subject category :

4. Rare Earthes and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Kowalczyk

Address for correspondence : Institute of Molecular Physics, PAS Smoluchowskiego 17, 60-179 Poznań Poland

Email address : ankow@ifmpan.poznan.pl